OpenStudy (anonymous):

If I have T to the 6 : (x,y) --> (x+2, y-1) how do I find the inverse?

OpenStudy (anonymous):

Find inverse of f(x)=(x+2)/(x-1). Step 1: replace f(x) with y: y = (x + 2)/(x - 1) Step 2: solve for x: (x - 1)y = x + 2 yx - y = x + 2 yx - x = y + 2 x(y - 1) = (y + 2) x = (y + 2)/(y - 1) Step 3: reverse roles of x & y: …

OpenStudy (anonymous):

Does this help?

OpenStudy (anonymous):

I'm not sure yet. Still thinking about it

OpenStudy (anonymous):

ok let me know if it does

OpenStudy (anonymous):

I guess I get that inverses reverse the x and y but I don't understand why this T is to the 6th power? is hat T( T(T(T(T(t(x)? Wouldn't I reverse this after taking the vector backward throug this?

OpenStudy (anonymous):

No you wouldn't

OpenStudy (anonymous):

can you tell me why? I get the inverse T and T^-1 but I don't understand what the t^6 indicates

OpenStudy (anonymous):

Do you know what T^-1 is right?